OpenStudy (anonymous):

The critical angle of refraction for calcite is 68.4 when it forms a boundary with water. Use this information to determine the speed of light in Calcite. Assume the refractive index of water ,1.333, is less than the refractive index of calcite.

OpenStudy (anonymous):

\[\color{green}{\dfrac{\sin \angle i_c}{\sin 90^{\circ}} = \dfrac{1.33}{n}}\] Solve for n.

OpenStudy (anonymous):

what is the original equation that we use ?

OpenStudy (anonymous):

Snell's Law: \[\color{green}{\dfrac{\sin \angle i_c}{\sin \angle \text{R}} = \dfrac{\text{n}_2}{\text{n}_1}}\]

OpenStudy (anonymous):

ok thanks :), How do you medal?

OpenStudy (anonymous):

what do I substitute for sin i then ?

OpenStudy (anonymous):

\[\color{green}{\dfrac{\sin 68.4^{\circ}}{\sin 90^{\circ}} = \dfrac{1.33}{n}}\]

OpenStudy (anonymous):

thank you